OpenStudy (anonymous):

find the derivative of f(x)= log5(3(x^4+5)^8)/tan(3x^2) (Do not simplify!) (Step by step) WILL FAN AND MEDAL!!

OpenStudy (anonymous):

\[f(x)=\log_{5}(3(x ^{4}+5)^{^{8}})\div \tan(3x ^{2}) \]

OpenStudy (anonymous):

\[the derivative of \log is \log_{b} u=du/u \ln b\] so the derivative of the log thing on top \[(24(x ^{4}+5)^{7}(4x ^{3}))/(3(x ^{4}+5)^{8}*\ln5\]

OpenStudy (anonymous):

derivative of tan(3x^2) is \[\sec ^{2}(3x ^{2})*(6x)\]

OpenStudy (anonymous):

do the quotient rule lowdhigh-highdlow all over low squared

OpenStudy (anonymous):

so its tan(3x^2)(derivative of the log thing)-(log thing)(derivative of tan) all over the tan thing squared